OpenStudy (anonymous):

Each box is packed with 40 units of components. An inspector will randomly select 5 units from a box to test if the components are in good condition. A box will be rejected if 2 or more tested units are defective. A shipment includes 400 boxes of the components. Each box will be tested by the inspector. The whole shipment will be returned if more than 90 boxes of components are rejected. The ordering records in the past showed that 20% of the boxes were rejected. What is the probability that a given shipment will be returned?

OpenStudy (anonymous):

Is there enough information given in this problem? My initial impression is no.

OpenStudy (anonymous):

But this seems like all the question has for me to answer it. And because this is an exam question, I think there should have enough information provided to solve the problem.

OpenStudy (zarkon):

looks binomial

OpenStudy (anonymous):

@Zarkon, also seems like hypergeomtric!

OpenStudy (zarkon):

but that part doesn't give you any information

OpenStudy (anonymous):

My initial intuition was binomial too. But if this is binomial, it means that the selection of the boxes for the inspection has replacement, which I don't think it has.

OpenStudy (anonymous):

hmm..but wouldn't there be a chance that one of the selected boxes for inspection is a repeated box? ie, a box that has already been inspected earlier?

OpenStudy (zarkon):

each box is being tested...the box is either good or bad...the probability it is bad is 20%

OpenStudy (anonymous):

Of course without replacement!

OpenStudy (anonymous):

20% 400 = 80

OpenStudy (anonymous):

...Doing it as binomial seems like a serious pain. You'd have to add up P(89) + P(88)+ P(87) +... + P(2) + P(1)

OpenStudy (zarkon):

you could use a normal approximation

OpenStudy (zarkon):

I'm all about the normal approximation today ;)

OpenStudy (anonymous):

Tell me more, Zarkon. How could we determine the SD?

OpenStudy (anonymous):

Could the numbers be used as a hypergeometric case? I was trying to figure out the numbers from this question for the the hypergeometric formula's parameters but doesn't look possible.

OpenStudy (anonymous):

haha...normal approximation today...

OpenStudy (zarkon):

if it is binomial with n trials and p prob of success...then the sd is \[\sqrt{n\cdot p\cdot (1-p)}\]

OpenStudy (zarkon):

brb

OpenStudy (anonymous):

Oh really? That's pretty cool. That does it for you then. You have the mean will clearly be 80. That SD gives you a z score for 90. Problem's doneso.

OpenStudy (anonymous):

Normal approximation was the first thing I thought of, but I figured you needed a given SD. It makes sense that you can calculate it though.

OpenStudy (anonymous):

This is assuming that there is replacement. If there is replacement, it will follow binomial distribution. However, there is a possibility that we picked out another box which we had tested earlier.

OpenStudy (anonymous):

I don't think so, EnOnn. It says that they simply test all 400 boxes. There's no chance of testing a box twice.

OpenStudy (zarkon):

you don't...you are looking at all the boxes...and checking each one once.

OpenStudy (zarkon):

all you need is independence between boxes

OpenStudy (anonymous):

ohhhh yea you are right! I was looking at the units! It is the units of components that are only randomly picked.

OpenStudy (anonymous):

Thanks a lot!! :D

OpenStudy (anonymous):

oh yes zarkon, in this case, they are all indepent. Thanks!! :D

OpenStudy (anonymous):

independent*

OpenStudy (anonymous):

Mean: 80 SD: sqrt(400*0.2*(.8)) = 8 90 is a z-score of 1.25

OpenStudy (anonymous):